OpenStudy (anonymous):

Last year the girls basketball team sold t=shirts and hoodies at the beginning of the season. The team sold a total of 42 items. Once the team won the 11th region tital they decided to sell t-shirts and hoodies again. They sold ten times as many t-shirts and 4 times as many hoodies as before. The number of items sold after the region equals 312. A. Assign labels to the unknowns B. Write two equations to form a system to solve for the unknowns C. Solve the systemm D. How many of each item were sold after the girls won the region title.

OpenStudy (anonymous):

So: Here we go: a) T = tshirts and H = hoodies. b) set up equations for the totals: T + H = 42 10T + 4H = 312 c) solve the system, by multiplying top equation by -4. (T + H = 42)* -4 10T + 4H = 312 ---------------------- -4T +-4 H = -168 10T + 4H = 312 (add equations) ------------------------ 6T = 144 T = 24. Then T + H = 42 -> 24 + H = 42, H = 18. d) then, since the after tourney is given by 10T + 4H = 312, put in 10*24 = 240 T-shirts, and 4*18 = 72 Hoodies, check by making sure: 72+240=312 --> and it does! 312 = 312, so the answer is right!

OpenStudy (anonymous):

Thank you you awesome. I was having a hard time. I have another.